w = EI the problem wants a solution for I, and W is wattage, however, it does not state if it is 110 watt or 220 watt, E is for voltaqge and I is for resistance. My question is how do I know what the the resistance is if I don't know what the voltage is, (110/220W)? Can u help?
Answer by ankor@dixie-net.com(15649)

:
w = E*I, where E = voltage and I = current (not resistance)
:
Perhaps they want you to assume 120 volts which is the voltage at a standard outlet.
:
Ohm's law is helpful:
Resistance: R = E/I,
Current: I = E/R,
Voltage: E = IR
:
More power (in watts) equations:
P = I^2 * R
P = E^2/R